OpenStudy (anonymous):

Which of the following is not an identity? A. cos^2 x csc x - csc x = -sin x B. sin x(cot x + tan x) = sec x C. cos^2 x - sin^2 x = 1- 2sin^2 x D. csc^2 x + sec^2 x = 1 ***i don't really get this one :( @terenzreignz? what do u think?

OpenStudy (anonymous):

ughh.. last one, and i don't get it :( *yells at brain* :P

terenzreignz (terenzreignz):

Which one does your instinct dictate? :D

OpenStudy (anonymous):

lol, well earlier, my instincts took me to the wrong place haha, so let's give it a try again this time... :p I think it's either A or C (again hahah looks like my instincts have quite the distinct pattern lol) but I'm leaning towards A... can you have a -sin x at the end of the equal sign? :/

terenzreignz (terenzreignz):

I see... I'll let @sriramkumar have his say... ;)

OpenStudy (anonymous):

this is the second chance i was talking about hahah question is, can the feelings/instinct redeem themselves?

OpenStudy (anonymous):

kk :)

OpenStudy (anonymous):

yep its the last once since we have csc^2x = 1/sin^2x and sec^2x = 1/cos^2x then you'll have it 1/cos^2x + 1/sin^2x = 1 cos^2x + sin^2x / cos^2x sin^2x = 1 1=cos^2x.sin^2x and its true for only some or none values of x, so it can't be an identity since an identity should be true for all values of x. hope u understood

terenzreignz (terenzreignz):

Proving something isn't an identity... well, if you could find a value of x which would make an equation false, then that equation is not an identity... Try having \(\huge x = \frac{\pi}{4}\) for letter D

OpenStudy (anonymous):

or simply say x=0... put it in all equations and see which one doesn't satisfy??

terenzreignz (terenzreignz):

That won't do it, @sriramkumar :P

terenzreignz (terenzreignz):

Or on second, thought, it will :D My bad...

OpenStudy (anonymous):

so you get 1=1 right?

OpenStudy (anonymous):

??? @terenzreignz an equation has a limited set of satisfying values, and an identity is true for all values of the x :P so try substituting 2 or 3 values of x then that would do

terenzreignz (terenzreignz):

Hardly... when \[\huge \sec\left(\frac{\pi}{4}\right)=\csc\left(\frac{\pi}{4}\right)=\sqrt2\]

OpenStudy (anonymous):

okay, so then you'd get 2=1 ? :/ cuz sq.rt. 2 + sq.rt. 2= sq.rt. 4 which equals 2 right? but if thats the case, then this isn't possible right? cuz 2 does not equal 1... :/ but maybe i did something wrong... which is quite likely :/

OpenStudy (anonymous):

so you want me to solve and show all the cards in there??? I can say that all the first three can be solved on for RHS but not in 4, you can't get RHS

OpenStudy (anonymous):

what's RHS? :/

OpenStudy (anonymous):

just Right Hand SIde :P :D :/

OpenStudy (anonymous):

haha ohh oaky i see :P

OpenStudy (anonymous):

still you'vent understood?

terenzreignz (terenzreignz):

In any case, @iheartfood , letting \[\large x = \frac{\pi}{4}\] should do it for choice D However... \[\Large \sqrt2 + \sqrt2 + \sqrt 4\] ???? Maybe you need more ice cream :P

terenzreignz (terenzreignz):

Besides... the csc and the sec were SQUARED in equation D.

OpenStudy (anonymous):

so since that's the case, D is the one that isn't an identity? cuz you're supposed to get 1=1 right? but here, we got 4=1? @terenzreignz does the thingy you put above get 4?

terenzreignz (terenzreignz):

Yes, it does get 4, @iheartfood And yes, D is the one which isn't an identity :P

OpenStudy (anonymous):

yay!! :) so in order for it to be an identity, it'd have to be 1=1 right?

terenzreignz (terenzreignz):

No, for it to be an identity you have to prove it, by turning the LHS into the RHS using some manipulations... And you can't do that in D.

OpenStudy (anonymous):

ahhh i see.. :) so as long as they equal each other in the end?

terenzreignz (terenzreignz):

Yes. Proving identities takes some creativity. Would you like me to demonstrate?

OpenStudy (anonymous):

yes please! :) an example would be great to help clear things up! :)

terenzreignz (terenzreignz):

Which one? A B and C are identities, so which one do you want me to prove...?

OpenStudy (anonymous):

can you please do A? :) and possibly C after if you have time? :)

terenzreignz (terenzreignz):

Well okay. First, let's start with the Pythagorean identity, you know this, right? \[\huge \sin^2(x) + \cos^2(x) = 1\]

OpenStudy (anonymous):

mhm :)

terenzreignz (terenzreignz):

Then it follows that \[\huge \sin^2(x)= 1-\cos^2(x)\] and \[\huge \cos^2(x) = 1-\sin^2(x)\] right?

OpenStudy (anonymous):

I'm following :)

terenzreignz (terenzreignz):

So from the first one, it also follows that \[\large -\sin^2(x)=-[1-\cos^2(x)]=\cos^2(x)=1\] catch me so far?

OpenStudy (anonymous):

yuppp :)

terenzreignz (terenzreignz):

rats... I meant \[\large -\sin^2(x)=-[1-\cos^2(x)]=\cos^2(x)-1\] \[\huge -\sin^2(x)=\cos^2(x)-1\]

terenzreignz (terenzreignz):

Okay?

OpenStudy (anonymous):

lol.... okay :)

terenzreignz (terenzreignz):

Now, to proceed with A. We start with the LHS (Left-Hand side) \[\large \cos^2(x)\csc(x) - \csc(x)\]

terenzreignz (terenzreignz):

Now, since \[\large \csc(x) = \frac1{\sin(x)}\] The LHS may be written as... \[\large \frac{\cos^2(x)}{\sin(x)}-\frac1{\sin(x)}\] Catch me so far?

OpenStudy (anonymous):

yeah making sense :)

terenzreignz (terenzreignz):

Now, since they have the same denominator, it can be written as \[\large \frac{\cos^2(x)-1}{\sin(x)}\]

OpenStudy (anonymous):

ahhh i see :)

terenzreignz (terenzreignz):

Now, the numerator can be simplified... into what? (refer to the preliminary identities I showed you :) )

OpenStudy (anonymous):

okay, hmmm tricky tricky! :P urmm do i still need to include the cos ? ;/ or does it change to something else?

terenzreignz (terenzreignz):

Scroll up, we derived an equation... using the Pythagorean identities... remember?

OpenStudy (anonymous):

lol its like you explain, and i can follow cuz its there! but then i have to fill in the blank parts, and its confusion and chaos! at least in my brain haha gotta work on this stuff ;/

OpenStudy (anonymous):

ohhh i see,,, so it simplifies to this? sin^2 x ?

terenzreignz (terenzreignz):

Are you sure? Check it again...

OpenStudy (anonymous):

urmm.... lemme see.. i got that from this part sin2(x)=1−cos2(x) ^^^ but does it not apply like that? where i can just flip the order around? :/

terenzreignz (terenzreignz):

I took a screen-cap... just to make a point :P

OpenStudy (anonymous):

ohh i see.. so it simplifies to -sin^2(x) ?

terenzreignz (terenzreignz):

Yes. For heaven's sake, these things I tell you, don't forget them :P \[\huge \frac{-\sin^2(x)}{\sin(x)}\] Which obviously simplifies into...?

OpenStudy (anonymous):

-sin(x) ?

terenzreignz (terenzreignz):

Yes. And as you can see, we have (magically) turned the LHS into the RHS (Right-hand side) Et voila I guess :P

OpenStudy (anonymous):

lol wow! i see :) thanks!!! :D

terenzreignz (terenzreignz):

Now, C.

OpenStudy (anonymous):

kk :)

terenzreignz (terenzreignz):

Start with the LHS as usual... \[\large \cos^2(x) - \sin^2(x) \] Now, we can express \(\cos^2(x)\) in terms of \(\sin^2(x)\) Once again, into what? Scroll up again, see the alternate way we can express \(\cos^2(x)\)

OpenStudy (anonymous):

this part right? −sin2(x)=−[1−cos2(x)]=cos2(x)=1

terenzreignz (terenzreignz):

No.

terenzreignz (terenzreignz):

You said you could follow :/

OpenStudy (anonymous):

yeah =( still a bit confused tho :/ it is also this right? 1−sin^2(x)

terenzreignz (terenzreignz):

Yes... again, to make a point :P (you DID say you were following...)

terenzreignz (terenzreignz):

Anyway... \[\Large \cos^2(x) - \sin^2(x) \]You can replace the \(\cos^2(x)\) with...?

OpenStudy (anonymous):

haha yeah i recall that :)

OpenStudy (anonymous):

just got a bit lost just now :) recovered! so i replace it with 1-sin^2 x ?

terenzreignz (terenzreignz):

Okay. \[\Large 1-\sin^2(x) - \sin^2(x)\] Simplifying yields...?

OpenStudy (anonymous):

1-2 sin^2 x ?

terenzreignz (terenzreignz):

Which is the RHS. Once again, we have managed to turn the LHS into the RHS... as if by magic :P This was basically an overview.. better consult your instructor on proving identities :) That was fun, no? :P

OpenStudy (anonymous):

haha yeah! its like a miracle! :P lol and yeah i will! thank you sooo much!! :) and lol super fun!! minus the parts where i experienced short-term memory and temporarily got lost hahaha :P yayy!! :) thank you so much for helping me @terenzreignz :)